OpenStudy (anonymous):

Ammonium bisulfide, NH4HS, forms ammonia, NH3, and hydrogen sulfide, H2S, through the reaction NH4HS(s)⇌NH3(g)+H2S(g) This reaction has a Kp value of 0.120 at 25 ∘C. What is the initial pressure of H2S(g) in the flask?

OpenStudy (anonymous):

0.065

OpenStudy (anonymous):

PV=nRT=mRT/M --> P=mRT/MV=0.45g*0.082(atm·l/K·mol)*298 K/ 34 (g/mol)*5 l= 0.065 atm